A corporate bond with $1000 maturity value carries a 7.5% coupon rate. It currently makes interest payments semi-annually. (i) This 12-year bond currently sells for $961.88. What is the rate of return on this bond? (ii) If the bond sold for $1,030.32, what is the rate of return on this bond?
1
Expert's answer
2016-02-22T00:01:00-0500
a) RM1000 maturity value, 7.5% coupon rate, interest payments semi-annually.
(i) If this 12-year bond currently sells for RM961.88, then the rate of return on this
bond is: Rate of return = (C + P1 - P0)/P0 = (1,000*0.075/2*24 + 961.88 - 1,000)/1,000 = 86.2%
(ii) If the bond sold for RM1,030.32, then the rate of return on this bond is: Rate of return = (C + P1 - P0)/P0 = (1,000*0.075/2*24 + 1,030.32 - 1,000)/1,000 = 93%